# Equilibrium Portfolio Selection under Utility-Variance Analysis of Log Returns in Incomplete Markets

### v1

@@ 2 Problem Formulation
    For any $k\geq 1$, there exists a constant $C(\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T,k,\|\pi\|_{\textup{BMO}})$ such that E_t[sup_s[t,T]—ξ_s^t,ε,π—^2k]C(r_,μ_,σ_,T,k,π_BMO)(ε—η—^2)^k, a.s. for $\varepsilon\in(0,T-t)$ sufficiently small.
@@
    For any $c>0,\pi\in\Pi_{0},\varepsilon\in(0,T-t)$, it holds that $\sup\limits_{\varepsilon\in(0,T-t)}\mathbb{E}_{t}\left[\exp(c|\xi^{t,\varepsilon,\pi}_{T}|)\right]<C(c,\pi,\|\eta\|_{\infty})<\infty,\mathbb{P}-a.s.$, where $C(c,\pi,\|\eta\|_{\infty})$ denotes a constant depending only on $c,\pi,\|\eta\|_{\infty}$, but independent of $\epsilon$, and is increasing with respect to $\|\eta\|_{\infty}$.
@@
The first assertion is standard, see Theorem 3.4.3 in zhang_backward_2017. The proof of the second assertion is similar to the ones of Lemma 2.4 and Lemma 2.5 in Hamaguchi_2021. Indeed, for any set $A\in\mathcal{F}_{t}$, it holds that
Here, we have omitted the dependence of $C$ for simplicity. Then, it follows that
$\mathbb{E}_{t}\left[\sup\limits_{s\in[t,T]}|\xi_{s}^{t,\varepsilon,\pi}|^{2k}\right]\leq C(\varepsilon|\eta|^{2})^{k}\left(1+\mathbb{E}_{t}\left[\left(\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)^{k}\right]\right)\leq C(\varepsilon|\eta|^{2})^{k},$
For the third assertion, note that
$\mathbb{E}_{t}\left[\exp(c|\xi^{t,\varepsilon,\pi}_{T}|)\right]$
$\leq$ $C(c,\eta)\left(\mathbb{E}_{t}\left[\exp\left(2c\|\sigma\|^{2}_{\infty}|\eta|\int_{t}^{T}|\pi_{s}|\mathrm{d}s\right)\right]\right)^{\frac{1}{2}}\left(\mathbb{E}_{t}\left[\exp\left(2c\int_{t}^{t+\varepsilon}\sigma_{s}\eta\mathrm{d}B_{s}-2c^{2}\int_{t}^{t+\varepsilon}|\sigma_{s}\eta|^{2}\right)\right]\right)^{\frac{1}{2}}$
$\leq$ $C(c,\delta,\|\eta\|_{\infty})\left(\mathbb{E}_{t}\left[\exp\left(\delta\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)\right]\right)^{\frac{1}{2}}\left(\mathbb{E}_{t}\left[\exp\left(2c\int_{t}^{t+\varepsilon}\sigma_{s}\eta\mathrm{d}B_{s}-2c^{2}\int_{t}^{t+\varepsilon}|\sigma_{s}\eta|^{2}\mathrm{d}s\right)\right]\right)^{\frac{1}{2}}.$
In view that $\sigma$ and $\eta$ are bounded, the Novikov condition is satisfied such that the second term equals $1$ a.s.. We also note that $C(c,\delta)$ can be chosen independent of $\varepsilon$. In fact, we can choose
$C(c,\delta,\|\eta\|_{\infty})=\exp\left(c(T-t)\left(\|\mu-r\|_{\infty}\cdot\|\eta\|_{\infty}+\frac{1}{2}\|\sigma\|_{\infty}^{2}\|\eta\|_{\infty}^{2}\right)+c^{2}(T-t)\|\sigma\|_{\infty}^{2}\|\eta\|_{\infty}^{2}+\frac{T}{\delta}\left(2c\|\sigma\|^{2}_{\infty}\|\eta\|_{\infty}\right)^{2}\right).$
As a result,
$\sup\limits_{\varepsilon\in(0,T-t)}\mathbb{E}_{t}\left[\exp(c|\xi^{t,\varepsilon,\pi}_{T}|)\right]\leq C(c,\delta,\|\eta\|_{\infty})\left(\frac{1}{1-\delta\|\pi\|_{\textup{BMO}}^{2}}\right)^{\frac{1}{2}}<\infty\quad a.s..$
@@
The above result is similar to Lemma 2.5 in Hamaguchi_2021. However, as the control process appears in the SDE governing the difference process, $\xi^{t,\varepsilon,\pi}$ cannot satisfy the properties in Hamaguchi_2021 under arbitrary trading strategies. This motivates us to consider the BMO condition into the definition of our admissible strategies. Consequently, the proof differs from that of Lemma 2.5 in Hamaguchi_2021.
@@
Assume that $\pi$ is an admissible trading strategy, for any $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$, $\pi^{t,\varepsilon,\eta}$ is an admissible strategy.
@@
We can choose $1<p^{\prime}<p$ and $q>1$ such that $\frac{1}{p}+\frac{1}{q}=\frac{1}{p^{\prime}}$. By Lemma 2.7, it holds that
$\leq$ $\left(\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]\right)^{\frac{1}{p}}\left(\mathbb{E}_{t}\left[\exp\left(q\zeta\left|\xi_{T}^{t,\varepsilon,\pi}\right|\right)\right]\right)^{\frac{1}{q}}<\infty\;a.s..$
Moreover, it is obvious that under a bounded perturbation $\eta$,
Thus, $\pi^{t,\varepsilon,\eta}$ is an admissible strategy. ∎
@@
For $\hat{\pi}\in\Pi_{0}$, $\hat{\pi}$ is called a time-consistent equilibrium strategy if
@@
In the following sections, we often fix a candidate equilibrium strategy $\hat{\pi}$ or an arbitrary admissible strategy $\pi$ for analysis. In proving the necessary conditions or verifying equilibrium, we will work with the perturbed log return process under a fixed perturbation $\eta$. To ease presentation, we denote by $\hat{\pi}^{t,\epsilon},\pi^{t,\epsilon}$ the perturbation strategy $\hat{\pi}^{t,\epsilon,\eta},\pi^{t,\epsilon,\eta}$, and denote by $\hat{R},R$ the log return process ${R}^{\hat{\pi}},R^{\pi}$, and denote by $\hat{R}^{t,\varepsilon},R^{t,\varepsilon}$ the perturbed log return process ${R}^{\hat{\pi}^{t,\epsilon,\eta}},{R}^{\pi^{t,\epsilon,\eta}}$, and denote by $\xi^{t,\varepsilon}$ the difference process $\xi^{t,\varepsilon,\pi}$ under the perturbation. For simplicity, we also use the notations $\hat{u}:=\sigma\hat{\pi}$ and $u:=\sigma\pi$ in the following sections.
@@ 3 A Necessary Condition of Time-Consistent Equilibria
In this section, we establish several properties of admissible strategies and characterize the necessary conditions of a time-consistent equilibrium by a BSDE system. Without loss of generality, we consider a fixed $p$ such that any admissible strategy satisfies
$\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]<\infty,\quad a.s..$
@@
$\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}.$
@@
The equilibrium strategy takes the form $\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}Z-\gamma\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma}$. We observe that the BSDE system admits a solution when $Z=0$ and $\tilde{Z}=0$, thus an equilibrium is given by
$\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta}{(\zeta+1)e^{-\zeta Y}+\gamma},$
where $Y$ is the solution of the ODE with terminal condition:
@@
Let $A=e^{-\zeta Y}$, then $A$ satisfies
$A_{t}=A(f(A)):=-\frac{\zeta(\zeta+1)A^{3}\theta^{2}}{2((\zeta+1)A+\gamma)^{2}}+\frac{\zeta A^{2}\theta^{2}}{(\zeta+1)A+\gamma}+r\zeta A,$ (3.4)
with the terminal condition $A(T)=1$. It is clear that $A$ has a lower bound 0. Combining $A^{\prime}=A(f(A))$ and the fact that $f(A)$ is bounded, we deduce that $A$ is bounded from above and thus there exists a unique solution and $A$ is positive.
@@
Hamaguchi_2021 derives a one-dimensional FBSDE as a necessary condition for an equilibrium in a general non-exponential time-inconsistent problem. In our setting, however, the introduced variance must to be represented by another BSDE of $\tilde{Y}$, which leads to strong coupling in our equation. Although dai_dynamic_2021 study a mean-variance framework including a variance term in their objective function, it is straightforward to see that the mean term can be represented by the BSDE of $\tilde{Y}$. In fact, $\tilde{Y}:=\mathbb{E}_{t}[R_{T}-R_{t}]$ (see in Lemma 3.4 later) corresponds precisely to the mean term at time $t$. Hence, their analysis also focus on a one-dimensional BSDE. Unlike Hamaguchi_2021 and dai_dynamic_2021, we focus on a two-dimensional fully coupled quadratic BSDE system.
To prove Theorem 3.1, we need some auxiliary preparations.
@@
$J(t,R_{t},\pi^{t,\varepsilon})-J(t,R_{t},\pi)$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(R_{s}+Y_{s}-R_{t})}(a(s,\pi^{t,\varepsilon}_{s})-a(s,\pi_{s})-\zeta(Z_{s}^{1}+\sigma_{s}\pi_{s}+\rho Z)\cdot\sigma_{s}\eta)\mathrm{d}s\right]$
@@
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma_{s}\pi_{s}+\tilde{Z}_{s}^{1}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}\pi_{s}^{t,\varepsilon}+\tilde{Z}_{s}^{1}+\rho\tilde{Z}_{s})^{2})\mathrm{d}s\right].$
Proof.
Fix the strategy $\pi$, $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$ and $\varepsilon\in(0,T-t)$. By the definition of $J$, we have
$J(t,R_{t},\pi^{t,\varepsilon})-J(t,R_{t},\pi)=\mathbb{E}_{t}\left[U(R_{T}^{t,\varepsilon}-R_{t})-U(R_{T}-R_{t})\right]-\frac{\gamma}{2}\left(\textup{Var}_{t}[R_{T}^{t,\varepsilon}]-\textup{Var}_{t}[R_{T}]\right)$
For the first term, we have that
$\mathbb{E}_{t}\left[U(R_{T}^{t,\varepsilon}-R_{t})-U(R_{T}-R_{t})\right]=e^{\zeta R_{t}}\mathbb{E}_{t}\left[U^{{}^{\prime}}(R_{T})\xi_{T}^{t,\varepsilon}+\int_{0}^{1}U^{{}^{\prime\prime}}(R_{T}+\lambda\xi_{T}^{t,\varepsilon})(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon}|^{2}\right].$
Next we analyze the conditional expectation $\mathbb{E}_{t}\left[U^{{}^{\prime}}(R_{T})\xi_{T}^{t,\varepsilon}\right]$ and let $(\alpha,\beta)$ be the unique adapted solution of the BSDE
Moreover, by taking the conditional expectations, we see that
[…]
$+\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(R_{s}+Y_{s})}\left[(\sigma_{s}\eta-\zeta\xi_{s}^{t,\varepsilon}(Z_{s}^{1}+\sigma_{s}\pi_{s}))\mathrm{d}B_{s}-\zeta\xi_{s}^{t,\varepsilon}Z_{s}\mathrm{d}\bar{B}_{s}\right]\right].$ (3.5)
Under Assumption 2.1, we have
The Burkholder-Davis-Gundy inequality yields that
For the second part, let $\tilde{\alpha}_{t}=\mathbb{E}_{t}[R_{T}]$ and $\tilde{Y_{t}}=\tilde{\alpha}_{t}-R_{t}$. Then, $\tilde{\alpha}$ and $\tilde{Y}$ satisfy that
and $\tilde{\alpha}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R})),\tilde{\beta}^{1}=\tilde{Z}^{1}+\sigma\pi,\tilde{\beta}=\tilde{Z}\in L^{2}_{\mathbb{F}}\left(\Omega;L^{2}(0,T;\mathbb{R})\right)$. From Lemma 2.4, it follows that $R^{\pi}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$ and hence $\tilde{Y}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$. We thus get that
$\textup{Var}_{t}[R_{T}]-\textup{Var}_{t}[R_{T}^{t,\varepsilon}]$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma\pi_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}-(\sigma\pi_{s}^{t,\varepsilon}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2})\mathrm{d}s\right],$ (3.6)
where we have used the fact that $R_{T}-R_{t+\varepsilon}=R_{T}^{t,\varepsilon}-R_{t+\varepsilon}^{t,\varepsilon}$ in the second and third equalities . The conclusion then follows by combining (3) and (3). ∎
@@
Let $\pi\in\Pi_{0}$ fixed. For any $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$ fixed and for $\varepsilon$ sufficiently small, it holds that a.s.
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon}|^{2}\right]\leq C(\|\eta\|_{\infty})|\eta|^{2},\quad a.s.$
with $C$ denoting a constant independent of $\varepsilon$ and increasing with respect to $\|\eta\|_{\infty}$.
Proof.
Therefore, we only need to show that $\frac{1}{\varepsilon}\mathbb{E}_{t}\left[e^{-(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}|\xi_{T}^{t,\varepsilon}|^{2}\right]\leq C|\eta|^{2}$ a.s. for $\lambda=0,1$, and for any $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$. By Lemma 2.7 and the Hölder’s inequality, there exists a constant $C$ only depending on $q,\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T,k,\|\pi\|_{\textup{BMO}}$ such that, for $\varepsilon$ sufficiently small,
$\leq C(\|\eta\|_{\infty})|\eta|^{2},\quad\lambda=0,1.$
@@
For any admissible strategy $\pi\in\Pi_{0}$, consider the pair (Y,Y~)($\tilde{Y}$) and (Z,Z~)($\tilde{Z}$) in Lemma 3.4. We deduce the existence of a measurable set $E_{1}\subset[0,T)$ with $\operatorname{Leb}([0,T]\setminus E_{1})=0$ such that, for any $t\in E_{1}$, there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}\subset(0,T-t)$ satisfying $\lim_{n\to\infty}\varepsilon_{n}^{t}=0$ and for any perturbation $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$, it holds that
@@
Proof.
$=$ $\alpha_{t}\left((\mu_{t}-r_{t})\eta-\frac{1}{2}\sigma^{2}_{t}\eta^{2}-\sigma^{2}_{t}\eta\pi_{t}-\zeta(Z_{t}^{1}+\sigma_{t}\pi_{t}+\rho Z_{t})\cdot\sigma_{t}\eta\right),\quad a.s.$
and
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon_{n}^{t}}2\sigma_{s}(\sigma_{s}\pi_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s}-\sigma_{s}^{2}\eta^{2})\mathrm{d}s\right]=2\sigma_{t}(\sigma_{t}\pi_{t}+\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t}-\sigma_{t}^{2}\eta^{2}),\quad a.s..$
@@
Note that $a(s,\pi^{t,\varepsilon_{n}^{t}}_{s})-a(s,\pi_{s})$ depends on the perturbation strategy $\pi^{t,\varepsilon_{n}^{t}}$ only on $t,t+\varepsilon_{n}^{t}]$ and will be affected by $\eta$, we cannot directly apply Lemma 3.3 in Hamaguchi_2021. However, because the effect of $\eta$ can be separated, we can apply Lemma 3.3 to each element in two expressions and then use the diagonal selection argument to complete the proof.
@@
Assume that $\hat{\pi}\in\Pi_{0}$ is an equilibrium strategy. Fix an arbitrary $\delta>0$ and take an arbitrary $t$ from the set $E_{1}$ in Lemma 3.6 . Let $\{\delta_{m}\}_{m\in\mathbb{N}}$ be a sequence such that $0<\delta_{m}\leq\delta,\;m\in\mathbb{N},$ and $\lim_{m\rightarrow\infty}\delta_{m}=0$. For each $m\in\mathbb{N}$, define $\eta_{m}$ by
@@
By the definition of the equilibrium, Lemma 3.6, Lemma 3.4 as well as Lemma 3.5, we obtain that there exists a sequence $\{\varepsilon_{n}^{t}\}_{n\in\mathbb{N}}\subset(0,T-t)$ satisfying $\lim_{n\to\infty}\varepsilon_{n}^{t}=0$ such that

### v2

@@ 2 Problem Formulation
    For any $k\geq 1$ and $\varepsilon\in(0,T-t)$, there exists a constant $C(k,\|\pi\|_{\textup{BMO}})$ [^6] Throughout the remainder of the paper, the constant $C$ may vary from line to line. For simplicity, we will omit its dependence on market parameters such as $\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T$, etc. Instead, we will explicitly indicate the dependence of $C$ on the trading strategy $\pi$ as well as on the perturbation parameters $\varepsilon$ and $\eta$ (see Definition 2.5) when they appear. such that E_t[sup_s[t,T]—ξ_s^t,ε,π—^2k]C(k,π_BMO)(ε—η—^2)^k, a.s..
@@
    For any $c\in\mathbb{R}$, it holds that $\sup\limits_{\varepsilon\in(0,T-t)}\mathbb{E}_{t}\left[\exp(c\xi^{t,\varepsilon,\pi}_{T})\right]\leq C(|c|,\|\pi\|_{\textup{BMO}},\|\eta\|_{\infty}),\,\mathbb{P}-a.s.$. Moreover, $C(|c|,\|\pi\|_{\textup{BMO}},\|\eta\|_{\infty})$ is non-decreasing with respect to $\|\eta\|_{\infty}$.
@@
See Appendix D.2. ∎
@@
The above result is similar to Lemma 2.5 in Hamaguchi_2021. However, the difference process $\xi^{t,\varepsilon,\pi}$ in (2.3) contains quadratic terms of the strategy $\pi\in\Pi$, and therefore does not satisfy the properties in Hamaguchi_2021 under arbitrary trading strategies. This motivates us to impose a BMO condition in the definition of admissible strategies. As a consequence, the proof differs from that of Lemma 2.5 in Hamaguchi_2021.
@@
Assume that $\pi\in\Pi$. Then for any $t\in[0,T)$, $\epsilon\in(0,T-t)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$, $\pi^{t,\varepsilon,\eta}\in\Pi$.
@@
See Appendix D.3. ∎
@@
For $\hat{\pi}\in\Pi$, $\hat{\pi}$ is called a time-consistent equilibrium strategy if
@@
In the following sections, we often fix a candidate equilibrium strategy $\hat{\pi}\in\Pi$ or an arbitrary strategy $\pi\in\Pi$ for analysis. In proving the necessary conditions or verifying theorems for equilibrium strategies, we will work with the perturbed log return processes ${R}^{\hat{\pi}^{t,\epsilon,\eta}}$ or ${R}^{\pi^{t,\epsilon,\eta}}$ under a fixed perturbation $\eta$. To ease presentation, we denote the perturbed strategies $\hat{\pi}^{t,\varepsilon,\eta}$ and $\pi^{t,\varepsilon,\eta}$ simply by $\hat{\pi}^{t,\varepsilon}$ and $\pi^{t,\varepsilon}$, respectively. Likewise, we write $\hat{R}$ and $R$ for the log-return processes $R^{\hat{\pi}}$ and $R^{\pi}$, and $\hat{R}^{t,\varepsilon}$ and $R^{t,\varepsilon}$ for the perturbed log-return processes $R^{\hat{\pi}^{t,\varepsilon}}$ and $R^{\pi^{t,\varepsilon}}$. The difference processes $\xi^{t,\varepsilon,\eta,\hat{\pi}}$ and $\xi^{t,\varepsilon,\eta,\pi}$ and are denoted simply by $\hat{\xi}^{t,\varepsilon}$ and $\xi^{t,\varepsilon}$. For simplicity, we also use the notations $\hat{u}:=\sigma\hat{\pi}$ and $u:=\sigma\pi$ in the following sections.
@@ 3 A Necessary Condition of Time-Consistent Equilibria
In this section, by establishing several properties of admissible strategies, we derive a necessary condition for a time-consistent equilibrium characterized by a coupled quadratic BSDE system. The main theorem of this section is stated below.
@@
$\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}.$ (3.2)
@@
The equilibrium strategy takes the form $\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}Z-\gamma\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma}$. We observe that the BSDE system admits a solution when $Z=0$ and $\tilde{Z}=0$, thus an equilibrium is given by $\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta}{(\zeta+1)e^{-\zeta Y}+\gamma}$, where $Y$ is the solution of the ODE with terminal condition:
@@
Let $A=e^{-\zeta Y}$, then $A$ satisfies the following ODE
$\left\{\begin{aligned} &A_{t}=A(f(A)),\quad t\in[0,T),\\ &A(T)=1,\end{aligned}\right$. (3.5)
where
$f(A):=-\frac{\zeta(\zeta+1)A^{2}\theta^{2}}{2((\zeta+1)A+\gamma)^{2}}+\frac{\zeta A\theta^{2}}{(\zeta+1)A+\gamma}+r\zeta.$
It is clear that $A$ has a lower bound 0. Combining $A^{\prime}=A(f(A))$ and the fact that $f(A)$ is bounded, we deduce that $A$ is bounded from above and thus there exists a unique solution and $A$ is positive.
@@
Hamaguchi_2021 derives a one-dimensional FBSDE as a necessary condition for an equilibrium in a general non-exponential discounting time-inconsistent problem. In our setting, however, the introduced variance must be represented by another BSDE for $\tilde{Y}$, which leads to strong coupling in our system. Although dai_dynamic_2021 study an MV framework that includes a variance term in the objective function, it is straightforward to see that the mean term can be represented by the BSDE for $\tilde{Y}$. In fact, $\tilde{Y}:=\mathbb{E}_{t}[R_{T}-R_{t}]$ (see Lemma 3.4 below) corresponds precisely to the mean term at time $t$. Hence, their analysis can also focus on a one-dimensional BSDE. Unlike Hamaguchi_2021 and dai_dynamic_2021, our problem naturally leads to a two-dimensional fully coupled quadratic BSDE system.
To prove Theorem 3.1, we first develop several auxiliary results. Lemma 3.4 characterizes the difference between the objective functional of a strategy $\pi\in\Pi_{p}$ and that of its perturbed counterpart. This lemma is inspired by horst2014forward, Hamaguchi_2021, and dai_dynamic_2021, and constitutes a key ingredient in establishing Theorem 3.1. The main idea in proving Lemma 3.4 is to apply a Taylor expansion and then compute the resulting expectations via a suitably constructed BSDE with an appropriate terminal condition. The proof of Lemma 3.4 is similar to that of Lemma 3.2 in Hamaguchi_2021. We include it in Appendix D.4 because our setting does not fit exactly within the framework of Hamaguchi_2021 or other existing results.
@@
$J(t,\pi^{t,\varepsilon})-J(t,\pi)$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(R_{s}+Y_{s}-R_{t})}(a(s,\pi^{t,\varepsilon}_{s})-a(s,\pi_{s})-\zeta(Z_{s}^{1}+\sigma_{s}\pi_{s}+\rho Z_{s})\cdot\sigma_{s}\eta)\mathrm{d}s\right]$
@@
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma_{s}\pi_{s}+\tilde{Z}_{s}^{1}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}\pi_{s}^{t,\varepsilon}+\tilde{Z}_{s}^{1}+\rho\tilde{Z}_{s})^{2})\mathrm{d}s\right].$ (3.6)
Lemma 3.4 shows that, for a fixed strategy $\pi\in\Pi_{p}$ ($p>1$), the equality (3.4) provides an explicit characterization of the difference between the objective functional of $\pi$ and that of its perturbed counterpart through the solution of the BSDE system (3.1) (with $\hat{\pi}$ replaced by $\pi$). In contrast, the following Corollary 3.5 provides the same characterization by using only a simplified form of (3.4), under the additional assumption that the simplified BSDE system (3.7) admits a solution. This corollary will be used in several subsequent verification theorems. The proof of Corollary 3.5 follows similar lines to that of Lemma 3.4, and a sketch of the proof is provided in Appendix D.5.
Corollary 3.5.
For any $p>1$ and $\pi\in\Pi_{p}$, assume that there exist continuous and adapted processes (Y,Y~)($\tilde{Y}$) such that $\exp\left(-\zeta(R+Y)\right)\in L^{p}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$ and $\tilde{Y}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$, as well as progressively measurable processes (Z,Z~)($\tilde{Z}$) such that $\int_{0}^{T}\left(|Z_{s}|^{2}+|\tilde{Z}_{s}|^{2}\right)\mathrm{d}s<\infty\,\,a.s..$ Moreover, suppose that (Y,Y~,Z,Z~)($\tilde{Y},$$\tilde{Z}$) satisfy the following BSDE
$\left\{\begin{aligned} &\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|\sigma_{s}\pi_{s}+\rho Z_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{s}|^{2}-a(s,\pi_{s})\right)\mathrm{d}s+Z_{s}\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}\tilde{Y}_{s}=-a(s,\pi_{s})\mathrm{d}s+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &Y_{T}=0,\\ &\tilde{Y_{T}}=0.\end{aligned}\right$. (3.7)
Then for any $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$ and $\varepsilon\in(0,T-t)$, we have, almost surely,
$J(t,\pi^{t,\varepsilon})-J(t,\pi)$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(R_{s}+Y_{s}-R_{t})}(a(s,\pi^{t,\varepsilon}_{s})-a(s,\pi_{s})-\zeta(\sigma_{s}\pi_{s}+\rho Z_{s})\cdot\sigma_{s}\eta)\mathrm{d}s\right]$
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma_{s}\pi_{s}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}\pi_{s}^{t,\varepsilon}+\rho\tilde{Z}_{s})^{2})\mathrm{d}s\right].$
To calculate $\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}(J(t,\pi^{t,\varepsilon})-J(t,\pi))$, we rely on Lemmas 3.6 and 3.7 below. Lemma 3.6 provides an upper bound for the second-order term appearing in the expansion of $J(t,\pi^{t,\varepsilon})-J(t,\pi)$ given in Lemma 3.4, while Lemma 3.7 establishes the convergence of the remaining two terms after division by $\varepsilon$. The proofs of Lemmas 3.6 and 3.7 are given in Appendix D.6 and Appendix D.7, respectively.
@@
For any $p>1$, $\pi\in\Pi_{p}$, $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$, it holds that
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon}|^{2}\right]\leq C(\|\eta\|_{\infty})|\eta|^{2},\quad a.s.,$
where $C$ is a constant independent of $\varepsilon$ and is non-decreasing with respect to $\|\eta\|_{\infty}$.
@@
For any $p>1$ and $\pi\in\Pi_{p}$, consider the pair (Y,Y~)($\tilde{Y}$) and $(Z,Z^{1},\tilde{Z},\tilde{Z}^{1})$ in Lemma 3.4. We deduce the existence of a measurable set $E_{1}\subset[0,T)$ with $\operatorname{Leb}([0,T]\setminus E_{1})=0$ such that, for any $t\in E_{1}$, there exists a sequence $\{\varepsilon_{n}^{t}\}_{n\in\mathbb{N}}\subset(0,T-t)$ satisfying $\lim\limits_{n\to\infty}\varepsilon_{n}^{t}=0$ and for any perturbation $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$, it holds that
@@
Note that the constant $C$ in Lemma 3.6 depends on $\|\eta\|_{\infty}$. Fortunately, $C$ is non-decreasing in $\|\eta\|_{\infty}$, which is sufficient for establishing the necessary condition; see the proof of Theorem 3.1. Moreover, since the integrands in (3.7) and (3.7) involve $a(s,\pi^{t,\varepsilon_{n}^{t}}_{s})$ and $(\sigma_{s}\pi_{s}^{t,\varepsilon}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}$, which depend on the perturbed strategy $\pi^{t,\varepsilon_{n}^{t}}$ over the interval $t,t+\varepsilon_{n}^{t}]$ and are thus influenced by the perturbation $\eta$, Lemma 3.3 in Hamaguchi_2021 cannot be applied directly. However, because the effect of $\eta$ can be separated, we may still use a diagonal selection argument to complete the proof, with $\varepsilon_{n}^{t}$ chosen independently of $\eta$.
@@
[…]
@@
By the definition of the equilibrium, Lemma 3.7, Lemma 3.4 as well as Lemma 3.6, we obtain that there exists a sequence $\{\varepsilon_{n}^{t}\}_{n\in\mathbb{N}}\subset(0,T-t)$ satisfying $\lim\limits_{n\to\infty}\varepsilon_{n}^{t}=0$ such that